# VisionReward: Fine-Grained Multi-Dimensional Human Preference Learning for Image and Video Generation

### v1

@@ 3.2 MPO for Text-to-Image Optimization
Dataset & Training Settings. The MPO training prompts are sampled from the VidProM [41] dataset. To adapt these prompts for video generation, we have optimized them following guidelines from CogVideoX [47], which results in roughly 22,000 samples. Due to limitations in the GPU budget, we have not used as many prompts as for image MPO, and we leave the data scaling study for future work. For our backbone model, we select CogVideoX-2b [^1] https://huggingface.co/THUDM/CogVideoX-2b. From each prompt, we have generated four videos. We use VisionReward to score these videos and apply the MPO algorithm to select approximately 9,400 effective preference pairs. In all our experiments, we maintain a batch size of 32, a learning rate of 5e-6, and employ 100 warmup steps followed by linear decay. We set the DPO parameter $\beta$ to 500. The MPO training process spans around 500 steps, equivalent to about 2 epochs. During training, we save a checkpoint every 40 steps and use a validation set split from the training set to pick the checkpoint with the highest reward. All experiments are conducted on 32 A800 GPUs.
@@
Evaluation Settings. To comprehensively assess the MPO models, we have conducted both automatic and human evaluations. The automatic evaluation is conducted across various benchmarks, including VBench [17] and our Video-MonetBench. For VBench, we focus on commonly reported key metrics, including Human Action, Scene, Multiple Objects, and Appearance Style (Appear. Style).
The human evaluation examines aspects like alignment to instructions, video realism, and aesthetic appeal. Annotators are asked to watch two videos and select the better one, with an option to indicate a tie if neither stands out. This manual assessment also uses the tasks in Video-MonetBench. In all these experiments, we utilize prompt optimization recommended in CogVideoX. Our baseline comparisons include the original CogVideoX-2b and VideoScore with MPO.
Figure 5: Human evaluation of text-to-video MPO.
@@
Experimental Results. The main results are shown in  Tab. 11 and  Tab. 12. When compared to the original CogVideoX-2b, MPO with VisionReward significantly enhances model performance across these benchmarks. In contrast, MPO with VideoScore tends to degrade performance on several metrics. This highlights the excellence of VisionReward in multi-objective reward modeling.
Human evaluation results, depicted in  Fig. 5, further validate the effectiveness of our approach. VisionReward with MPO demonstrates significant improvements over the base model, achieving approximately a 27% increase in win rate. Additionally, VisionReward with MPO largely outperforms the VideoScore with MPO, underscoring the superiority of VisionReward compared to VideoScore.
Ablation Study: Different Strategy for MPO. In using the MPO algorithm, we define the way in which $R^{i}$ dominates $R^{j}$ (given two images $x^{i}$ and $x^{j}$) to select pairs. The definition of “dominate” includes at least three methods for different reward objective: total weighted score, score of each dimension, and score of each sub-dimension. To investigate the impact of different definitions of “dominate”, we conduct experiments based on CogVideoX-5b. Specifically, we employ three different strategies, setting the threshold of the total score to 0.8, 0.6, and 0.4 respectively, ensuring that the number of pairs obtained through all three strategies is 5k. We use a batch size of 64, a learning rate of 2e-6, the DPO parameter $\beta$ of 500, and the training steps of 300. After training, we compare the evaluation results of VisionReward.  Tab. 13 shows that using the score for each dimension as the reward objective yields the best results.
Methods Baseline Total Dimension Sub-dimension
VisionReward 4.303 4.515 4.573 4.514
Table 13: Score of VisionReward after different strategies of MPO. Total: “dominate” based on total weighted score. Dimension: “dominate” based on score of each dimension. Sub-dimension: “dominate” based on score of each sub-dimension.
@@ 5 Conclusion
We introduce VisionReward, a reward model for visual generation, which is fine-grained and multi-dimensional. By enabling Vision-Language Model (VLM) to perform checklist assessments and applying linear summation with weighting coefficients derived from preference learning, VisionReward achieves highly accurate and interpretable. Based on VisionReward, we develop an algorithm of multi-objective preference learning, which addresses issues of over-optimization or lack-optimization.
@@ 1.1 Details of Annotation Pipeline
Details of Data Preparation. To ensure the diversity of our annotation, we collect our annotated data from various sources. For images, we select 3 popular datasets: ImageRewardDB [46], HPDv2 [44], and Pick-a-Pic [20]. We sample 16k images from each dataset for annotation (with the first two sampling 4k prompts each associated with 4 images, and Pick-a-Pic sampling 8k prompts each associated with 2 images), totaling 48k annotated images. For videos, we conduct random sampling from the VidProM dataset and use ChatGLM [10] for data cleaning (prompt Cf.  Tab. 14), leading to 10k prompts. Then we use CogVideoX [47], VideoCrafter2 [4] and OpenSora [51] to generate 30k videos, sample from Panda-70M [5] to get 3k real videos, leading to 33k videos for annotation.
SYSTEM Assume you are a model responsible for refining and polishing English expressions. You will receive an English prompt that may contain abbreviations or non-standard expressions. Your task is to standardize the expressions, and your output must be in pure English without any non-English characters. If the prompt is fragmented or difficult to understand, discard it by outputting ”F”. Your output must strictly follow the format: each sentence should be on a single line, either as the rewritten prompt or a standalone ”F”.
@@
Annotation Design. In the annotation process, we design a comprehensive dimensional system and divide each dimension into several sub-dimensions. For each sub-dimension, we set options that vary gradually in degree. Annotators are required to label each sub-dimension for every image.  Tab. 15 and  Tab. 16 show the annotation taxonomy of images, while  Tab. 17 and  Tab. 18 show for videos. After the annotation is completed, we map the options of each sub-dimension into checklist questions, which are then used for model training.
@@ 1.2 Statistics of Annotation Result
(a) Text-to-image
(b) Text-to-video
Figure 6: Annotation statistics of different sub-dimensions.
@@
2 More Analysis on VisionReward
Algorithm 2 Iterative Regression with Weight Masking
2:  Initialize convergence criterion $\text{diff}\leftarrow\infty$
3:  while $\text{diff}>\epsilon$ do
4:     $_{\text{old}}\leftarrow\textbf{w}$ w
5:     for each $(\mathbf{X}_{i},\mathbf{X}_{j},y)$ in $\mathcal{D}$ do
6:        $\Delta\mathbf{X}\leftarrow\mathbf{X}_{i}-\mathbf{X}_{j}$
7:        $\hat{y}\leftarrow\sigma(\Delta\mathbf{X}^{T}\mathbf{w})$
8:        $\nabla_{\textbf{w}}\text{Loss}=(\hat{y}-y)\Delta\mathbf{X}$
9:        $\mathbf{w}\leftarrow\mathbf{w}-\alpha\nabla_{\textbf{w}}\text{Loss}$
11:     $\text{diff}\leftarrow||\textbf{w}-\textbf{w}_{\text{old}}||$
11:  Trained weights $\mathbf{w}$
@@ 2 More Analysis on VisionReward
Weight Masking. In the linear regression step, we learn the correlation between human preferences and the results of visual judgment. In our design, if the result of the visual judgment is ”yes”, human preference improves. We examine the correlation between human preference and each judgment result, and the numerical results indicate a positive correlation. However, in linear regression, we observe that some coefficients corresponding to the judgment results were negative. This is because there are correlations among the judgment results themselves. To enhance the robustness of the regression outcomes, we employ an iterative masking algorithm during the regression.
@@ 3 More Details of MonetBench
Details of Classification Proportions. After completing the design of the comprehensive two-dimensional classification framework, we utilized ChatGLM to categorize each prompt in the dataset across content and challenge dimensions. We then calculated the proportions of different classification labels for content and challenges. The content and challenge categories and their respective examples are summarized in Tables 23 and 24. Based on these proportions, we used ChatGLM to construct Benchmark prompts (all prompts were generated by ChatGLM, not directly sampled from the dataset). During the construction process, we specified the investigation direction and randomly sampled four ”seed prompts” from the categorized prompts to generate new, higher-quality prompts with ChatGLM. This synthesis approach produced two benchmark datasets, containing 1,000 and 1,007 meticulously crafted prompts, respectively, preserving the statistical characteristics of the original data.
The final datasets provide balanced and comprehensive coverage of content and challenge categories, as detailed in Tables 23 and 24. Table 25 lists the specific content and challenge categories with detailed descriptions and example prompts, providing a clear understanding of the dataset’s composition. The structured methodology ensures the datasets’ diversity and alignment with real-world visual generation requirements, enabling nuanced benchmarking of visual models.
@@
4 More Details and Results of MPO
4.1 Details of Prompt for MPO
  • Semantic-Based Filtering: Utilizing an existing scene graph parser [43], we evaluate the semantic richness of prompts by analyzing the number of subjective and objective relationships. Prompts with fewer than one meaningful relationship are filtered out to reduce noise in the dataset.
4.2 More Results of MPO
(a) Overall Score
(b) Composition Score
(c) Fidelity Score
(d) Alignment Score
(e) Quality Score
(f) Safety & Emotion Score
Figure 8: Qualitative result of MPO in text-to-image.
Figure 9: Qualitative result of MPO in text-to-video.
Figure 10: Qualitative result of MPO in text-to-video.

### v2

@@ 3.2 VisionReward for Text-to-Vision Optimization
Dataset & Training Settings. For our backbone model, we select CogVideoX-2B [^1] https://huggingface.co/THUDM/CogVideoX-2b. The training prompts are sampled from VidProM [44]. To adapt these prompts for video generation, we have optimized them following guidelines from CogVideoX [50], which results in roughly 22,000 samples. We generate 4 videos for each prompt, and use VisionReward to score these videos and apply the MPO strategy to select approximately 9,400 effective preference pairs. In all our experiments, we maintain a batch size of 32, a learning rate of 5e-6, and employ 100 warmup steps followed by linear decay. We set the DPO parameter $\beta$ to 500. The MPO training process spans around 500 steps, equivalent to about 2 epochs. During training, we save a checkpoint every 40 steps and use a validation set split from the training set to pick the checkpoint with the highest reward.
@@
Evaluation Settings. To comprehensively assess the MPO models, we have conducted both automatic and human evaluations. The automatic evaluation is conducted across various benchmarks, including VBench [19] and our Video-MonetBench. For VBench, we focus on commonly reported key metrics, including Human Action, Scene, Multiple Objects, and Appearance Style (Appear. Style). For human evaluation, annotators are asked to watch two videos and select the better one, with an option to indicate a tie if neither stands out. This manual assessment also uses the tasks in MonetBench. In all these experiments, we utilize prompt optimization recommended in CogVideoX. Our baseline comparisons include the original CogVideoX-2B and DPO with VideoScore.
@@
Experimental Results. The main results are shown in  Tab. 7 and  Tab. 8. When compared to the original CogVideoX-2B, optimization with VisionReward significantly enhances model performance across these benchmarks. In contrast, optimization with VideoScore tends to degrade performance. The empirical evidence substantiates VisionReward’s advanced capacity for multi-dimensional optimization. (Case study in Appendix  Sec. 9.3.)
3.3 Effects of Multi-Dimensional Optimization
Ablation Study: Different Strategy for MPO. In using the MPO strategy, we define the way in which $R^{i}$ dominates $R^{j}$ (given two images $x^{i}$ and $x^{j}$) to select pairs. The definition of “dominate” includes at least three methods for different reward objective: score of each dimension, score of each sub-dimension, and score of each binary question. To investigate the impact of different definitions of “dominate”, we conduct experiments based on CogVideoX-5B. Specifically, we employ three different strategies, setting the threshold of the total score to 0.6, 0.5, and 0.4 respectively, ensuring that the number of pairs obtained through all three strategies is 5k. We use a batch size of 64, a learning rate of 2e-6, the DPO parameter $\beta$ of 500, and the training steps of 300. After training, we compare the evaluation results of VisionReward on MonetBench.  Tab. 9 shows that using the score for each dimension as objective yields the best results.
Methods Baseline Dimension Sub-dimension Question
Reward 4.303 4.573 4.539 4.514
Table 9: Score of VisionReward after different strategies of MPO. Dimension: “dominate” based on score of each dimension. Sub-dimension: “dominate” based on score of each sub-dimension. Question: “dominate” based on score of each binary question.
Method Align. Quality Dynamic Physics Preserv. Overall
Original 1.733 0.660 0.053 0.344 0.653 4.303
DPO 1.697 0.680 0.034 0.356 0.741 4.515
DPO w/ MPO 1.766 0.688 0.042 0.356 0.721 4.573
MaPO 1.736 0.660 0.052 0.345 0.645 4.295
MaPO w/ MPO 1.737 0.656 0.055 0.349 0.649 4.321
Table 10: Ablation Study for MPO strategy. Scores are given by VisionReward on MonetBench.
Ablation Study: Multi-Dimensional or Single Score. To comprehensively illustrate how MPO addresses the factor optimization bias inherent in DPO, we extend the previously described experimental setup by utilizing the total score of VisionReward to select pairs, setting the threshold at 0.8 to ensure consistency in numbers. VisionReward is employed to evaluate scores across various dimensions, with the detailed results presented in  Tab. 10. Through the implementation of the MPO strategy, CogVideoX is optimized in such a manner that it avoids the degradation of certain factors (e.g., alignment), thereby achieving improved trade-offs, such as maintaining good preservation while avoiding excessively slow dynamic changes. The empirical evidence further substantiates VisionReward’s capacity for algorithm-agnostic preference alignment, as evidenced by comparative testing with other approaches like MaPO [16].
For efficiency discussion, the experimental results in  Tab. 11 reveal the comparative effectiveness of the MPO and DPO methods. These results suggest that MPO outperforms the DPO approach in terms of both efficiency and effectiveness.
These findings highlight promising directions for future research in developing novel optimization algorithms and adapting VisionReward for multi-dimensional optimization and preference balancing.
Method Threshold #Sample Overall
DPO 0.6 6473 4.442
DPO 0.8 5270 4.515
MPO 0.6 5028 4.573
Table 11: More comparison of MPO and DPO on MonetBench.
@@ 5 Conclusion
We introduce VisionReward, a reward model for visual generation, which is fine-grained and multi-dimensional. By enabling Vision-Language Model (VLM) to perform binary assessments and applying linear summation with weighting coefficients derived from preference learning, VisionReward achieves highly accurate and interpretable. For visual generative optimization, VisionReward surpasses other reward models and enable multi-dimensional strategy.
@@ 7.1 Details of Annotation Pipeline
Details of Data Preparation. To ensure the diversity of our annotation, we collect our annotated data from various sources, as presented in  Tab. 12. For images, we select 3 popular datasets: ImageRewardDB [49], HPDv2 [47], and Pick-a-Pic [22]. We sample 16k images from each dataset for annotation (with the first two sampling 4k prompts each associated with 4 images, and Pick-a-Pic sampling 8k prompts each associated with 2 images), totaling 48k annotated images. For videos, we conduct random sampling from the VidProM dataset and use ChatGLM [11] for data cleaning (prompt Cf.  Tab. 13), leading to 10k prompts. Then we use CogVideoX [50], VideoCrafter2 [4] and OpenSora [54] to generate 30k videos, sample from Panda-70M [5] to get 3k real videos, leading to 33k videos for annotation.
Type Source #Samples #Checklist
Image ImageRewardDB [49] 16K 1M
Pick-a-Pic [22] 16K 1M
HPDv2 [47] 16K 1M
Video CogVideoX [50] 10K 0.6M
Open-Sora [54] 10K 0.6M
VideoCrafter2 [4] 10K 0.6M
Panda-70M [5] 3K 0.2M
Table 12: Statistics of source data and annotation.
(a) Text-to-image
(b) Text-to-video
Figure 6: Annotation statistics of different sub-dimensions.
Annotation Design. In the annotation process, we design a comprehensive dimensional system and divide each dimension into several sub-dimensions.  Tab. 14 illustrates our preference dimensions and the checklist count. We identify 5 dimensions for text-to-image generation and expand to 9 dimensions for text-to-video. For each sub-dimension, we set options that vary gradually in degree. Annotators are required to label each sub-dimension for every image. In our annotation process, we design multiple-choice questions for each dimension, allowing annotators to select the option that best fits the image. Additionally, we further decompose these options into several binary (yes/no) questions, ultimately forming a checklist for evaluating each image or video. This checklist is then utilized for subsequent evaluation systems and reward model training.
Annotation Management. To avoid annotation bias of annotators, our approach includes:
  • Professional management. Cooperated with a specialized company, we strictly conduct trial annotation training for the annotators, screen qualified annotators, and conduct secondary quality inspection of the annotation results.
  • Standard document. Our annotation document gives clear definitions and provides more than 10 examples for each judgment, to align the standard among annotators.
  • Consistency test. Consistency of annotators on checklist results in 89.29% (images) and 89.33% (videos).
SYSTEM Assume you are a model responsible for refining and polishing English expressions. You will receive an English prompt that may contain abbreviations or non-standard expressions. Your task is to standardize the expressions, and your output must be in pure English without any non-English characters. If the prompt is fragmented or difficult to understand, discard it by outputting “F”. Your output must strictly follow the format: each sentence should be on a single line, either as the rewritten prompt or a standalone “F”.
@@
Dimension #Sub-dimension #Checklist
Image Video Image Video
Alignment 1 1 1 4
Composition 5 1 13 2
Quality 5 4 14 14
Fidelity 5 3 25 9
Safety&Emotion 2 1 8 4
Stability - 5 - 12
Dynamic - 2 - 8
Physics - 1 - 4
Preservation - 2 - 7
Total 18 20 61 64
Table 14: Taxonomy of annotation for VisionReward.
@@ 7.2 Statistics of Annotation Result
8 More Details and Results of VisionReward
Weight Masking. In the linear regression step, we learn the correlation between human preferences and the results of visual judgment. In our design, if the result of the visual judgment is “yes”, human preference improves. We examine the correlation between human preference and each judgment result, and the numerical results indicate a positive correlation. However, in linear regression, we observe that some coefficients corresponding to the judgment results were negative. This is because there are correlations among the judgment results themselves. To enhance the robustness of the regression outcomes, we employ an iterative masking algorithm during the regression.
Algorithm 2 Iterative Regression with Weight Masking
2:  Initialize convergence criterion $\text{diff}\leftarrow\infty$
3:  while $\text{diff}>\epsilon$ do
4:     $_{\text{old}}\leftarrow\textbf{w}$ w
5:     for each $(\mathbf{X}_{i},\mathbf{X}_{j},y)$ in $\mathcal{D}$ do
6:        $\Delta\mathbf{X}\leftarrow\mathbf{X}_{i}-\mathbf{X}_{j}$
7:        $\hat{y}\leftarrow\sigma(\Delta\mathbf{X}^{T}\mathbf{w})$
8:        $\nabla_{\textbf{w}}\text{Loss}=(\hat{y}-y)\Delta\mathbf{X}$
9:        $\mathbf{w}\leftarrow\mathbf{w}-\alpha\nabla_{\textbf{w}}\text{Loss}$
11:     $\text{diff}\leftarrow||\textbf{w}-\textbf{w}_{\text{old}}||$
11:  Trained weights $\mathbf{w}$
Ablation Study: Efficiency of VisionReward via Weight Masking. We explore whether selective question pruning through coefficient thresholding could maintain model efficacy. As demonstrated in  Tab. 15, coefficient-guided masking strategy (Acc. w/ mask) achieves remarkable efficiency gains with comparable accuracy.
#question 2 4 8 16 32 all
Acc. w/ mask 71.1 78.9 80.7 81.4 81.6 81.7
Table 15: Accuracy on HPDv2-test for different question scales. “Acc. w/ mask” means selecting questions by coefficient magnitude after full-set regression.
9 More Details and Results of MPO
9.1 Details of Prompt for MPO
  • Semantic-Based Filtering: Utilizing an existing scene graph parser [46], we evaluate the semantic richness of prompts by analyzing the number of subjective and objective relationships. Prompts with fewer than one meaningful relationship are filtered out to reduce noise data.
9.2 VisionReward for Text-to-Image Optimization
We maintain consistent training parameters and dataset sizes across all experiments to ensure fair comparison. For all three experiments, we used an effective batch size of 256 (with GAS set to 4 and train batch size set to 1), set $\beta$ to 5000, and a learning rate of 5e-9 (before scaling). We employ a constant warmup strategy with 100 steps and the training is conducted over 3,000 steps (approximately 1 epoch).
Experimental Results. Main results are demonstrated in  Tab. 16 and  Tab. 17. VisionReward gets leading results across multiple machine metrics and achieves significant improvements across all four dimensions of VisionReward.
Methods CLIP Aes HPSv2 PickScore
Original 0.273 5.463 0.282 22.25
Pick-a-Pic 0.279 5.511 0.286 22.45
HPSv2 0.277 5.599 0.292 22.58
VisionReward 0.279 5.612 0.289 22.61
Table 16: Evaluation results of multiple metrics on DrawBench.
Methods Composition Quality Fidelity Safety
Original 0.755 0.550 0.009 -0.008
Pick-a-Pic 0.765 0.588 0.009 -0.009
HPSv2 0.874 0.630 0.010 -0.004
VisionReward 0.894 0.670 0.017 -0.001
Table 17: Evaluation results on MonetBench.
9.3 More Results of MPO
Figure 7: Qualitative result of MPO in text-to-image.
Figure 8: Qualitative result of MPO in text-to-video.
Figure 9: Qualitative result of MPO in text-to-video.
(a) Overall Score
(b) Composition Score
(c) Fidelity Score
(d) Alignment Score
(e) Quality Score
(f) Safety & Emotion Score
10 Details of Dimensions for Annotation
Tab. 18 and  Tab. 19 show the annotation taxonomy of images, while  Tab. 20 and  Tab. 21 show for videos.
Correlation of Sub Dimensions. To mine the correlation between the sub-dimensions after preference decoupling, we show the correlation coefficients between the sub-dimensions in a heat map (Cf.  Fig. 11).
@@ 10 Details of Dimensions for Annotation
Figure 11: Correlation heatmap of text-to-video sub dimensions.
@@ 11 More Details of MonetBench
Details of Classification Proportions. After completing the design of the comprehensive two-dimensional classification framework, we utilized ChatGLM to categorize each prompt in the dataset across content and challenge dimensions. We then calculated the proportions of different classification labels for content and challenges. The content and challenge categories and their respective examples are summarized in Tables 26 and 27. Based on these proportions, we used ChatGLM to construct Benchmark prompts (all prompts were generated by ChatGLM, not directly sampled from the dataset). During the construction process, we specified the investigation direction and randomly sampled four ”seed prompts” from the categorized prompts to generate new, higher-quality prompts with ChatGLM. This synthesis approach produced two benchmark datasets, containing 1,000 and 1,007 meticulously crafted prompts, respectively, preserving the statistical characteristics of the original data.
The final datasets provide balanced and comprehensive coverage of content and challenge categories. Table 28 lists the specific content and challenge categories with detailed descriptions and example prompts, providing a clear understanding of the dataset’s composition. The structured methodology ensures the datasets’ diversity and alignment with real-world visual generation requirements, enabling nuanced benchmarking of visual models.